OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 6x – 3y = 12 and equation of the line passing through the point (–4, –2) and perpendicular to y = –x + 6 Thank you!

OpenStudy (anonymous):

@Lotsohelpneeded first thing u must know is how to find a slope of line can u find?????????

OpenStudy (anonymous):

I know how but it's hard for me because I have dyslexia

OpenStudy (anonymous):

1) Identify the correct slope and y intercept of the equation 6x – 3y = 12. Slope-Intercept form is: y = mx + b m is slope b is y-intercept SO, change it into Slope-Intercept form.~ 6x – 3y = 12 Subtract 6x from both sides. -3y = -6x + 12 Divide both sides by -3. y = 2x - 4 y = 2x - 4 Slope is 2 Y-intercept is -4

OpenStudy (anonymous):

Thanks jazy!

OpenStudy (anonymous):

Just one lest question? Please?

OpenStudy (anonymous):

*last

OpenStudy (anonymous):

You're welcome! I have to go now but maybe @Aperogalics can help you with the 2nd one. Oh and welcome to OpenStudy!

OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of –6 and a y-intercept of 9

OpenStudy (anonymous):

Thanks again

OpenStudy (anonymous):

Remember Slope-Intercept form is: y = mx + b m is slope b is y-intercept And plug in the values.(: y = -6x + 9

OpenStudy (anonymous):

@Lotsohelpneeded y=mx+c is the slope-intercept form that's not fair @jazy ...lol:)

OpenStudy (anonymous):

But wait can I still have the answer to the second problem I originally needed help with

OpenStudy (anonymous):

Jazy?

OpenStudy (anonymous):

Okay I ended up not leaving! I'll help with the original one.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

The equation of the line passing through the point (–4, –2) and perpendicular to y = –x + 6 The slope of a perpendicular line is opposite reciprocal. So if our given line has a slope of -1 or -1/1, the new line will have a slope of 1/1 or just 1. So you have: Slope of 1 Point (-4, -2) Use Point-Slope form. y - y1 = m(x - x1) y - (-2) = 1(x - (-2)) Two negatives (-) equal to a Positive(+) y + 2 = 1(x + 2) You need Slope-Intercept form. So, distribute and solve for y. y + 2 = 1(x + 2) y + 2 = x + 2 Subtract 2 from both sides. y = x + 0 Or y = x

OpenStudy (anonymous):

Thanks jazy!

OpenStudy (anonymous):

Welcome! (: And now I really do have to go. lol

OpenStudy (anonymous):

But I don;t have that on the answers

OpenStudy (anonymous):

they are only y=-x+6 y=x-6 y=-x-2 y=x+2

OpenStudy (anonymous):

nvm